please solve this one...
A man divides $10,000 in two investments, one at 10% and the other at 30%. Find how much is invested at each rate so that the two investments produce the same income annually.. the answer is $2,500 and $7,500.. I just need the solution thank you..

Answer by mananth(16949)

Investment I 10.00% per annum ---x
Investment II 30.00% per annum ---y

x + y= 10000 ------------------------1
10.00% x + 30.00% y = $860.00
Multiply by 100
10 x -30 y = 0 --------2

Multiply (1) by -10
we get

-10 x -10 y = -100000.00

Add this to (2)

0 x -40 y = -100000

divide by -40

y = 2500 investment at 30.00%
Balance 7500 investment at 10.00%